# Arcade Processes for Informed Martingale Interpolation

### v2

@@ 4.1 The one-arc FAM
$\,\mathrm{d}F^{X_{1}\,|\,X_{0},I_{t}^{(1)}=z}(y)=\frac{f^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(z)\,\mathrm{d}F^{X_{1}\,|\,X_{0}}(y)}{\int_{\mathbb{R}}f^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(z)\,\mathrm{d}F^{X_{1}\,|\,X_{0}}(y)}.$ (4.15)
Inserting the expression for $\,\mathrm{d}F^{X_{1}\,|\,X_{0},I_{t}^{(1)}=z}(y)$ into $M_{t}$, we obtain
$M_{t}=\int_{\mathbb{R}}y\,\mathrm{d}F^{X_{1}\,|\,X_{0},I_{t}^{(1)}}(y)=\frac{\int_{\mathbb{R}}yf^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(I_{t}^{(1)})\,\mathrm{d}F^{X_{1}\,|\,X_{0}}(y)}{\int_{\mathbb{R}}f^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(I_{t}^{(1)})\,\mathrm{d}F^{X_{1}\,|\,X_{0}}(y)}.$ (4.16)
Example 4.6.
$X_{1}\,|\,X_{0}=\left\{\begin{array}[]{ll}\frac{3}{2}X_{0}+\frac{1}{2}&\mbox{with probability }\frac{3}{4},\\ \\ -\frac{1}{2}X_{0}-\frac{3}{2}&\mbox{with probability }\frac{1}{4}.\end{array}\right.$ (4.17)
This is a martingale coupling since $\mathbb{E}[X_{1}\,|\,X_{0}]=X_{0}$. In fact, it can be shown that this coupling is the solution to a martingale optimal transport problem, see [25]. For any $\{T_{0},T_{1}\}$-conditionally-Markov $X$-RAP $(I_{t}^{(1)})$, we have
$M_{t}=\mathbb{E}[X_{1}\,|\,X_{0},I_{t}^{(1)}]=\frac{(9X_{0}+3)f^{I_{t}^{(1)}\,|\,X_{0},X_{1}=\frac{3}{2}X_{0}+\frac{1}{2}}(I_{t}^{(1)})-(X_{0}+3)f^{I_{t}^{(1)}\,|\,X_{0},X_{1}=-\frac{1}{2}X_{0}-\frac{3}{2}}(I_{t}^{(1)})}{6f^{I_{t}^{(1)}\,|\,X_{0},X_{1}=\frac{3}{2}X_{0}+\frac{1}{2}}(I_{t}^{(1)})+2f^{I_{t}^{(1)}\,|\,X_{0},X_{1}=-\frac{1}{2}X_{0}-\frac{3}{2}}(I_{t}^{(1)})}.$ (4.18)
Example 4.7.
Let $X_{0}\sim\mathcal{N}(0,1)$ and $X_{1}\sim\mathcal{N}(0,2)$, where $X_{1}\,|\,X_{0}\sim\mathcal{N}(X_{0},1)$. For any $\{T_{0},T_{1}\}$-conditionally-Markov $X$-RAP $(I_{t}^{(1)})$, we have
$M_{t}=\mathbb{E}[X_{1}\,|\,X_{0},I_{t}^{(1)}]=\frac{\int_{\mathbb{R}}yf^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(I_{t}^{(1)})\mathrm{e}^{\frac{-(y-X_{0})^{2}}{2}}\,\mathrm{d}y}{\int_{\mathbb{R}}f^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(I_{t}^{(1)})\mathrm{e}^{\frac{-(y-X_{0})^{2}}{2}}\,\mathrm{d}y}.$ (4.19)
@@
$u(t,z,X_{0},y):=f^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(z),\quad u_{t}(t,z,X_{0},y):=\frac{\partial}{\partial t}u(t,z,X_{0},y),$ (4.20)
$u_{z}(t,z,X_{0},y):=\frac{\partial}{\partial z}u(t,z,X_{0},y),\quad u_{zz}(t,z,X_{0},y):=\frac{\partial^{2}}{\partial z^{2}}u(t,z,X_{0},y),$ (4.21)
$K_{\boldsymbol{\cdot}}(t,z,X_{0}):=\int_{\mathbb{R}}u_{\boldsymbol{\cdot}}(t,z,X_{0},y)\,\mathrm{d}F^{X_{1}\,|\,X_{0}}(y),\quad V_{\boldsymbol{\cdot}}(t,z,X_{0}):=\int_{\mathbb{R}}yu_{\boldsymbol{\cdot}}(t,z,X_{0},y)\,\mathrm{d}F^{X_{1}\,|\,X_{0}}(y).$ (4.22)
@@
Let $M_{t}=\mathbb{E}[X_{1}\,|\,X_{0},I_{t}^{(1)}]$ be a one-arc FAM. If the $X$-RAP $(I_{t}^{(1)})$ is a semimartingale such that $(t,x)\rightarrow\frac{V(t,x,X_{0})}{K(t,x,X_{0})}$ is $C^{2}(((T_{0},T_{1})\setminus N)\times Im(I^{(1)}))$ where $N\subset(T_{0},T_{1})$ contains finitely many elements, then
@@
The Itô condition, i.e., $(t,x)\rightarrow\frac{V(t,x,X_{0})}{K(t,x,X_{0})}$ is $C^{2}(((T_{0},T_{1})\setminus N)\times Im(I^{(1)})))$, where $N\subset(T_{0},T_{1})$ contains finitely many elements, imposes implicit integrability conditions on $(X_{0},X_{1})$.
Since Gaussian RAPs have been studied in detail above, we can specialize Proposition 4.8 to this particular subclass.
@@
Under the conditions in Proposition 4.8, if the conditional probability distribution of $I_{t}^{(1)}$ given $(X_{0},X_{1})$ is $\mathcal{N}(g_{0}(t)X_{0}+g_{1}(t)X_{1}+\mu_{A}(t),\sigma_{A}^{2}(t))$, we have
$\frac{V_{t}(t,I_{t}^{(1)},X_{0})-M_{t}K_{t}(t,I_{t}^{(1)},X_{0})}{K(t,I_{t}^{(1)},X_{0})}=-\frac{(\mathbb{E}[X_{1}^{3}\,|\,X_{0},I_{t}^{(1)}]-M_{t}^{3})g_{1}(t)\left(\frac{g_{1}(t)}{\sigma_{A}(t)}\right)^{\prime}}{\sigma_{A}(t)}$
$\hskip 56.9055pt+\Bigg{(}\frac{M_{t}g_{1}(t)\left(\frac{g_{1}(t)}{\sigma_{A}(t)}\right)^{\prime}+(I_{t}^{(1)}-g_{0}(t)X_{0}-\mu_{A}(t))\left(\left(\frac{g_{1}(t)}{\sigma_{A}(t)}\right)^{\prime}+g_{1}(t)\left(\frac{1}{\sigma_{A}(t)}\right)^{\prime}\right)}{\sigma_{A}(t)}$
$\hskip 199.16928pt-\frac{(X_{0}g_{0}^{\prime}(t)+\mu_{A}^{\prime}(t))g_{1}(t)}{\sigma_{A}^{2}(t)}\Bigg{)}\times\mathrm{Var}[X_{1}\,|\,X_{0},I_{t}^{(1)}],$ (4.25)
$\frac{V_{z}(t,I_{t}^{(1)},X_{0})-M_{t}K_{z}(t,I_{t}^{(1)},X_{0})}{K(t,I_{t}^{(1)},X_{0})}=\frac{g_{1}(t)}{\sigma_{A}^{2}(t)}\mathrm{Var}[X_{1}\,|\,X_{0},I_{t}^{(1)}],$ (4.26)
$\frac{M_{t}K_{z}^{2}(t,I_{t}^{(1)},X_{0})-K_{z}(t,I_{t}^{(1)},X_{0})V_{z}(t,I_{t}^{(1)},X_{0})}{K^{2}(t,I_{t}^{(1)},X_{0})}=\frac{I_{t}^{(1)}-g_{0}(t)X_{0}-\mu_{A}(t)-g_{1}(t)M_{t}}{\sigma_{A}^{4}(t)}g_{1}(t)$
×$\hskip 341.43306pt\times\mathrm{Var}[X_{1}\,|\,X_{0},I_{t}^{(1)}],$ (4.27)
$\frac{V_{zz}(t,I_{t}^{(1)},X_{0})-M_{t}K_{zz}(t,I_{t}^{(1)},X_{0})}{2K(t,I_{t}^{(1)},X_{0})}=\frac{\mathbb{E}[X_{1}^{3}\,|\,X_{0},I_{t}]-M_{t}^{3}}{2\sigma_{A}^{4}(t)}g_{1}^{2}(t)$
$\hskip 170.71652pt-\frac{2(I_{t}^{(1)}-g_{0}(t)X_{0}-\mu_{A}(t))+g_{1}(t)M_{t}}{2\sigma_{A}^{4}(t)}g_{1}(t)\mathrm{Var}[X_{1}\,|\,X_{0},I_{t}^{(1)}].$ (4.28)
@@
Let $Z_{t}:=I_{t}^{(1)}-g_{0}(t)X_{0}-\mu_{A}(t)$ and $J_{t}:=X_{0}\,g_{0}^{\prime}(t)+\mu_{A}^{\prime}(t)$. The following computations of the partial derivatives $V_{t},K_{t},V_{z},V_{zz},K_{zz}$ in the case where $u(t,z,X_{0},y)=(1/\sqrt{2\pi\sigma_{A}^{2}(t)})\\ \exp[-(z-g_{0}(t)X_{0}-g_{1}(t)y-\mu_{A}(t))^{2}/(2\sigma^{2}_{A}(t))]$ lead to the result:
@@
    $V_{t}(t,I_{t}^{(1)},X_{0})$ $=\frac{g_{1}(t)\sigma_{A}^{\prime}(t)-g_{1}^{\prime}(t)\sigma_{A}(t)}{\sigma_{A}^{3}(t)}g_{1}(t)K(t,I_{t}^{(1)},X_{0})\mathbb{E}[X_{1}^{3}\,|\,X_{0},I_{t}^{(1)}]$
    $\quad+\frac{Z_{t}J_{t}\sigma_{A}(t)+\sigma_{A}^{\prime}(t)(Z_{t}^{2}-\sigma_{A}^{2}(t))}{\sigma_{A}^{3}(t)}V(t,I_{t}^{(1)},X_{0})$
    $\quad-\frac{g_{1}(t)(J_{t}\sigma_{A}(t)+2Z_{t}\sigma_{A}^{\prime}(t))-Z_{t}\sigma_{A}(t)g_{1}^{\prime}(t)}{\sigma_{A}^{3}(t)}K(t,I_{t}^{(1)},X_{0})\mathbb{E}[X_{1}^{2}\,|\,X_{0},I_{t}^{(1)}],$ (4.29)
@@
    $K_{t}(t,I_{t}^{(1)},X_{0})$ $=\frac{g_{1}(t)\sigma_{A}^{\prime}(t)-g_{1}^{\prime}(t)\sigma_{A}(t)}{\sigma_{A}^{3}(t)}g_{1}(t)K(t,I_{t}^{(1)},X_{0})\mathbb{E}[X_{1}^{2}\,|\,X_{0},I_{t}^{(1)}]$
@@
    $V_{z}(t,I_{t}^{(1)},X_{0})=\frac{g_{1}(t)\mathbb{E}[X_{1}^{2}\,|\,X_{0},I_{t}^{(1)}]K(t,I_{t}^{(1)},X_{0})-Z_{t}V(t,I_{t}^{(1)},X_{0})}{\sigma_{A}^{2}(t)},$ (4.31)
@@
    $V_{zz}(t,I_{t}^{(1)},X_{0})$ $=\frac{g_{1}^{2}(t)E[X_{1}^{3}\,|\,X_{0},I_{t}^{(1)}]-2Z_{t}g_{1}(t)\mathbb{E}[X_{1}^{2}\,|\,X_{0},I_{t}^{(1)}]}{\sigma_{A}^{4}(t)}K(t,I_{t}^{(1)},X_{0})$
@@
    $K_{zz}(t,I_{t}^{(1)},X_{0})$ $=\frac{g_{1}^{2}(t)E[X_{1}^{2}\,|\,X_{0},I_{t}^{(1)}]K(t,I_{t}^{(1)},X_{0})-2Z_{t}g_{1}(t)V(t,I_{t}^{(1)},X_{0})}{\sigma_{A}^{4}(t)}K(t,I_{t}^{(1)},X_{0})$
    $\quad+\frac{Z_{t}^{2}-\sigma_{A}^{2}(t)}{\sigma_{A}^{4}(t)}K(t,I_{t}^{(1)},X_{0}).$ (4.34)
Keeping the notation $Z_{t}=I_{t}^{(1)}-g_{0}(t)X_{0}-\mu_{A}(t)$ and $J_{t}=X_{0}\,g_{0}^{\prime}(t)+\mu_{A}^{\prime}(t)$ introduced in the proof, and introducing $U_{t}=\mathbb{E}[X_{1}^{3}\,|\,X_{0},I_{t}]-M_{t}^{3}$, the SDE satisfied by $(M_{t})$ can be rewritten as
$\,\mathrm{d}M_{t}=$ $U_{t}\left(\frac{g_{1}^{2}(t)}{2\sigma_{A}^{4}(t)}\,\mathrm{d}[I^{(1)}]_{t}-\frac{g_{1}(t)\left(\frac{g_{1}(t)}{\sigma_{A}(t)}\right)^{\prime}}{\sigma_{A}(t)}\,\mathrm{d}t\right)+\frac{g_{1}(t)}{\sigma_{A}^{2}(t)}\mathrm{Var}[X_{1}\,|\,X_{0},I_{t}^{(1)}]$
×$\times\Bigg{(}\Bigg{(}M_{t}\sigma_{A}(t)\left(\frac{g_{1}(t)}{\sigma_{A}(t)}\right)^{\prime}+Z_{t}\frac{\sigma_{A}(t)}{g_{1}(t)}\left(\left(\frac{g_{1}(t)}{\sigma_{A}(t)}\right)^{\prime}+g_{1}(t)\left(\frac{1}{\sigma_{A}(t)}\right)^{\prime}\right)-J_{t}\Bigg{)}\,\mathrm{d}t$
@@
If, furthermore, $(I_{t}^{(1)})$ is a standard RAP (see Definition 3.6), its driver $(D_{t})$ is Gauss-Markov with $K_{D}(x,y)=H_{1}(\min(x,y))H_{2}(\max(x,y))$, where $H_{1}$ and $H_{2}$ are continuous functions on $[T_{0},T_{1}]$ such that $H_{1}/H_{2}$ is positive and non-decreasing on $[T_{0},T_{1}]$. Then, as shown in [14],
$[I^{(1)}]_{t}=[D]_{t}=\int_{T_{0}}^{t}H_{2}(s)\,\mathrm{d}H_{1}(s)-\int_{T_{0}}^{t}H_{1}(s)\,\mathrm{d}H_{2}(s),$ (4.36)
@@
and $H_{1}/H_{2}$ is continuous and monotone, so of bounded variation and differentiable almost everywhere.
Recalling that in the standard RAP case we have
@@
$\,\mathrm{d}M_{t}=\frac{\mathrm{Var}[X_{1}\,|\,X_{0},I_{t}^{(1)}]}{H_{1}(T_{1})H_{2}(t)-H_{1}(t)H_{2}(T_{1})}$
×$\times\left(\left(\frac{Z_{t}(H_{1}^{\prime}(t)H_{2}(T_{1})-H_{1}(T_{1})H_{2}^{\prime}(t))-M_{t}(H_{1}^{\prime}(t)H_{2}(t)-H_{1}(t)H_{2}^{\prime}(t))}{H_{1}(T_{1})H_{2}(t)-H_{1}(t)H_{2}(T_{1})}-J_{t}\right)\,\mathrm{d}t+\,\mathrm{d}I^{(1)}_{t}\right),$ (4.38)
@@
$\,\mathrm{d}N_{t}:=\left(\frac{Z_{t}(H_{1}^{\prime}(t)H_{2}(T_{1})-H_{1}(T_{1})H_{2}^{\prime}(t))-M_{t}(H_{1}^{\prime}(t)H_{2}(t)-H_{1}(t)H_{2}^{\prime}(t))}{H_{1}(T_{1})H_{2}(t)-H_{1}(t)H_{2}(T_{1})}-J_{t}\right)\,\mathrm{d}t+\,\mathrm{d}I^{(1)}_{t},$ (4.41)
@@
$M_{t}=X_{0}+\int_{T_{0}}^{t}\frac{\mathrm{Var}[X_{1}\,|\,X_{0},I_{t}^{(1)}]}{H_{1}(T_{1})H_{2}(t)-H_{1}(t)H_{2}(T_{1})}\,\mathrm{d}N_{t}.$ (4.42)
As shown below, the process $(N_{t})_{t\in[T_{0},T_{1}]}$, defined by the SDE 4.41 and initial condition $N_{T_{0}}=0$, is a martingale, and can be used to construct a standard Brownian motion on $[T_{0},T_{1}]$.
@@
Let $I\subseteq\mathbb{R}^{+}$ be an interval or the real positive line. A stochastic process $(Y_{t})_{t\in I}$ is said to be standard Brownian motion on an interval $[a,b]\subseteq I$ if there exists a standard Brownian motion $(B_{t})_{t\geqslant 0}$ such that $Y_{t}\overset{a.s.}{=}B_{t-a}$ for all $t\in[a,b]$.
@@
Under the conditions in Proposition 4.8, the process $(W_{t})_{t\in[T_{0},T_{1}]}$ defined by
$W_{t}=\int_{T_{0}}^{t}\frac{1}{H_{1}^{\prime}(s)H_{2}(s)-H_{1}(s)H_{2}^{\prime}(s)}\,\mathrm{d}N_{s}$ (4.43)
is an $(\mathcal{F}_{t}^{I})$-adapted standard Brownian motion on $[T_{0},T_{1}]$.

### v3

@@ 5.1 The one-arc FAM
$\mathrm{d}F^{X_{1}\mid X_{0},I_{t}^{(1)}=z}(y)=\frac{f^{I_{t}^{(1)}\mid X_{0},X_{1}=y}(z)\,\mathrm{d}F^{X_{1}\mid X_{0}}(y)}{\int_{\mathbb{R}}f^{I_{t}^{(1)}\mid X_{0},X_{1}=y}(z)\,\mathrm{d}F^{X_{1}\mid X_{0}}(y)}.$ (5.17)
Inserting the expression for $\mathrm{d}F^{X_{1}\mid X_{0},I_{t}^{(1)}=z}(y)$ into $M_{t}$, we obtain
$M_{t}=\int_{\mathbb{R}}y\,\mathrm{d}F^{X_{1}\mid X_{0},I_{t}^{(1)}}(y)=\frac{\int_{\mathbb{R}}yf^{I_{t}^{(1)}\mid X_{0},X_{1}=y}(I_{t}^{(1)})\,\mathrm{d}F^{X_{1}\mid X_{0}}(y)}{\int_{\mathbb{R}}f^{I_{t}^{(1)}\mid X_{0},X_{1}=y}(I_{t}^{(1)})\,\mathrm{d}F^{X_{1}\mid X_{0}}(y)}.$ (5.18)
@@
$u(t,z,X_{0},y)\mathrel{\mathop{\ordinarycolon}}=f^{I_{t}^{(1)}\mid X_{0},X_{1}=y}(z),\quad u_{t}(t,z,X_{0},y)\mathrel{\mathop{\ordinarycolon}}=\frac{\partial}{\partial t}u(t,z,X_{0},y),$ (5.19)
$u_{z}(t,z,X_{0},y)\mathrel{\mathop{\ordinarycolon}}=\frac{\partial}{\partial z}u(t,z,X_{0},y),\quad u_{zz}(t,z,X_{0},y)\mathrel{\mathop{\ordinarycolon}}=\frac{\partial^{2}}{\partial z^{2}}u(t,z,X_{0},y),$ (5.20)
$K_{\boldsymbol{\cdot}}(t,z,X_{0})\mathrel{\mathop{\ordinarycolon}}=\int_{\mathbb{R}}u_{\boldsymbol{\cdot}}(t,z,X_{0},y)\,\mathrm{d}F^{X_{1}\mid X_{0}}(y),$ (5.21)
$V_{\boldsymbol{\cdot}}(t,z,X_{0})\mathrel{\mathop{\ordinarycolon}}=\int_{\mathbb{R}}yu_{\boldsymbol{\cdot}}(t,z,X_{0},y)\,\mathrm{d}F^{X_{1}\mid X_{0}}(y).$ (5.22)
@@
Several examples of one-arc FAMs are provided in B. Next, we we go about deriving the stochastic differential equation for a class of one-arc FAMs.
@@
Let $M_{t}=\mathbb{E}[X_{1}\mid X_{0},I_{t}^{(1)}]$ be a one-arc FAM. If the $X$-RAP $(I_{t}^{(1)})$ is a semimartingale such that $(t,x)\rightarrow\frac{V(t,x,X_{0})}{K(t,x,X_{0})}$ is $C^{2}(((T_{0},T_{1})\setminus N)\times Im(I^{(1)}))$ where $N\subset(T_{0},T_{1})$ contains finitely many elements, then
@@
The condition $(t,x)\rightarrow\frac{V(t,x,X_{0})}{K(t,x,X_{0})}$ is $C^{2}(((T_{0},T_{1})\setminus N)\times Im(I^{(1)})))$, where $N\subset(T_{0},T_{1})$ contains finitely many elements, imposes implicit integrability conditions on $(X_{0},X_{1})$.
Since Gaussian RAPs have been studied in detail in previous sections, we specialise Proposition 5.32 to this particular subclass.
@@
Under the conditions in Proposition 5.32, if the conditional probability distribution of $I_{t}^{(1)}$ given $(X_{0},X_{1})$ is $\mathcal{N}(g_{0}(t)X_{0}+g_{1}(t)X_{1}+\mu_{A}(t),\sigma_{A}^{2}(t))$ for all $t\in(T_{0},T_{1})$, we have
$\frac{V_{t}(t,I_{t}^{(1)},X_{0})-M_{t}K_{t}(t,I_{t}^{(1)},X_{0})}{K(t,I_{t}^{(1)},X_{0})}=-\frac{(\mathbb{E}[X_{1}^{3}\mid X_{0},I_{t}^{(1)}]-M_{t}^{3})g_{1}(t)\left(\frac{g_{1}(t)}{\sigma_{A}(t)}\right)^{\prime}}{\sigma_{A}(t)}$
$\hskip 28.45274pt+\Bigg(\frac{M_{t}g_{1}(t)\left(\frac{g_{1}(t)}{\sigma_{A}(t)}\right)^{\prime}+(I_{t}^{(1)}-g_{0}(t)X_{0}-\mu_{A}(t))\left(\left(\frac{g_{1}(t)}{\sigma_{A}(t)}\right)^{\prime}+g_{1}(t)\left(\frac{1}{\sigma_{A}(t)}\right)^{\prime}\right)}{\sigma_{A}(t)}$
$\hskip 170.71652pt-\frac{(X_{0}g_{0}^{\prime}(t)+\mu_{A}^{\prime}(t))g_{1}(t)}{\sigma_{A}^{2}(t)}\Bigg)\times\mathrm{Var}[X_{1}\mid X_{0},I_{t}^{(1)}],$ (5.25)
$\frac{V_{z}(t,I_{t}^{(1)},X_{0})-M_{t}K_{z}(t,I_{t}^{(1)},X_{0})}{K(t,I_{t}^{(1)},X_{0})}=\frac{g_{1}(t)}{\sigma_{A}^{2}(t)}\mathrm{Var}[X_{1}\mid X_{0},I_{t}^{(1)}],$ (5.26)
$\frac{M_{t}K_{z}^{2}(t,I_{t}^{(1)},X_{0})-K_{z}(t,I_{t}^{(1)},X_{0})V_{z}(t,I_{t}^{(1)},X_{0})}{K^{2}(t,I_{t}^{(1)},X_{0})}$
$\hskip 113.81102pt=\frac{I_{t}^{(1)}-g_{0}(t)X_{0}-\mu_{A}(t)-g_{1}(t)M_{t}}{\sigma_{A}^{4}(t)}g_{1}(t)\mathrm{Var}[X_{1}\mid X_{0},I_{t}^{(1)}],$ (5.27)
$\frac{V_{zz}(t,I_{t}^{(1)},X_{0})-M_{t}K_{zz}(t,I_{t}^{(1)},X_{0})}{2K(t,I_{t}^{(1)},X_{0})}=\frac{\mathbb{E}[X_{1}^{3}\mid X_{0},I_{t}]-M_{t}^{3}}{2\sigma_{A}^{4}(t)}g_{1}^{2}(t)$
$\hskip 113.81102pt-\frac{2(I_{t}^{(1)}-g_{0}(t)X_{0}-\mu_{A}(t))+g_{1}(t)M_{t}}{2\sigma_{A}^{4}(t)}g_{1}(t)\mathrm{Var}[X_{1}\mid X_{0},I_{t}^{(1)}].$ (5.28)
@@
Let $Z_{t}\mathrel{\mathop{\ordinarycolon}}=I_{t}^{(1)}-g_{0}(t)X_{0}-\mu_{A}(t)$ and $J_{t}\mathrel{\mathop{\ordinarycolon}}=X_{0}\,g_{0}^{\prime}(t)+\mu_{A}^{\prime}(t)$. The following computations of the partial derivatives $V_{t},K_{t},V_{z},V_{zz},K_{zz}$ in the case where $u(t,z,X_{0},y)=(1/\sqrt{2\pi\sigma_{A}^{2}(t)})\\ \exp[-(z-g_{0}(t)X_{0}-g_{1}(t)y-\mu_{A}(t))^{2}/(2\sigma^{2}_{A}(t))]$ lead to the result:
@@
    $V_{t}(t,I_{t}^{(1)},X_{0})=\frac{g_{1}(t)\sigma_{A}^{\prime}(t)-g_{1}^{\prime}(t)\sigma_{A}(t)}{\sigma_{A}^{3}(t)}g_{1}(t)K(t,I_{t}^{(1)},X_{0})\mathbb{E}[X_{1}^{3}\mid X_{0},I_{t}^{(1)}]$
    $\hskip 142.26378pt+\frac{Z_{t}J_{t}\sigma_{A}(t)+\sigma_{A}^{\prime}(t)(Z_{t}^{2}-\sigma_{A}^{2}(t))}{\sigma_{A}^{3}(t)}V(t,I_{t}^{(1)},X_{0})$
    $\quad-\frac{g_{1}(t)(J_{t}\sigma_{A}(t)+2Z_{t}\sigma_{A}^{\prime}(t))-Z_{t}\sigma_{A}(t)g_{1}^{\prime}(t)}{\sigma_{A}^{3}(t)}K(t,I_{t}^{(1)},X_{0})\mathbb{E}[X_{1}^{2}\mid X_{0},I_{t}^{(1)}],$ (5.29)
@@
    $K_{t}(t,I_{t}^{(1)},X_{0})$ $=\frac{g_{1}(t)\sigma_{A}^{\prime}(t)-g_{1}^{\prime}(t)\sigma_{A}(t)}{\sigma_{A}^{3}(t)}g_{1}(t)K(t,I_{t}^{(1)},X_{0})\mathbb{E}[X_{1}^{2}\mid X_{0},I_{t}^{(1)}]$
@@
    $V_{z}(t,I_{t}^{(1)},X_{0})=\frac{g_{1}(t)\mathbb{E}[X_{1}^{2}\mid X_{0},I_{t}^{(1)}]K(t,I_{t}^{(1)},X_{0})-Z_{t}V(t,I_{t}^{(1)},X_{0})}{\sigma_{A}^{2}(t)},$ (5.31)
@@
    $V_{zz}(t,I_{t}^{(1)},X_{0})$ $=\frac{g_{1}^{2}(t)E[X_{1}^{3}\mid X_{0},I_{t}^{(1)}]-2Z_{t}g_{1}(t)\mathbb{E}[X_{1}^{2}\mid X_{0},I_{t}^{(1)}]}{\sigma_{A}^{4}(t)}K(t,I_{t}^{(1)},X_{0})$
@@
    $K_{zz}(t,I_{t}^{(1)},X_{0})$ $=\frac{g_{1}^{2}(t)E[X_{1}^{2}\mid X_{0},I_{t}^{(1)}]K(t,I_{t}^{(1)},X_{0})-2Z_{t}g_{1}(t)V(t,I_{t}^{(1)},X_{0})}{\sigma_{A}^{4}(t)}$
    $\quad\times K(t,I_{t}^{(1)},X_{0})+\frac{Z_{t}^{2}-\sigma_{A}^{2}(t)}{\sigma_{A}^{4}(t)}K(t,I_{t}^{(1)},X_{0}).$ (5.34)
Keeping the notation $Z_{t}=I_{t}^{(1)}-g_{0}(t)X_{0}-\mu_{A}(t)$ and $J_{t}=X_{0}\,g_{0}^{\prime}(t)+\mu_{A}^{\prime}(t)$ introduced in the proof, and introducing $U_{t}=\mathbb{E}[X_{1}^{3}\mid X_{0},I_{t}]-M_{t}^{3}$, the SDE satisfied by $(M_{t})$ can be rewritten as
$\,\mathrm{d}M_{t}=$ $U_{t}\left(\frac{g_{1}^{2}(t)}{2\sigma_{A}^{4}(t)}\,\mathrm{d}[I^{(1)}]_{t}-\frac{g_{1}(t)\left(\frac{g_{1}(t)}{\sigma_{A}(t)}\right)^{\prime}}{\sigma_{A}(t)}\,\mathrm{d}t\right)+\frac{g_{1}(t)}{\sigma_{A}^{2}(t)}\mathrm{Var}[X_{1}\mid X_{0},I_{t}^{(1)}]$
$\times\Bigg(\Bigg(M_{t}\sigma_{A}(t)\left(\frac{g_{1}(t)}{\sigma_{A}(t)}\right)^{\prime}+Z_{t}\frac{\sigma_{A}(t)}{g_{1}(t)}\left(\left(\frac{g_{1}(t)}{\sigma_{A}(t)}\right)^{\prime}+g_{1}(t)\left(\frac{1}{\sigma_{A}(t)}\right)^{\prime}\right)-J_{t}\Bigg)\,\mathrm{d}t$
@@
If, furthermore, $(I_{t}^{(1)})$ is a standard RAP (see Definition 3.20), its driver $(D_{t})$ is Gauss-Markov with $K_{D}(x,y)=H_{1}(\min(x,y))H_{2}(\max(x,y))$, where $H_{1}$ and $H_{2}$ are continuous functions on $T_{0},T_{1}]$ such that $H_{1}/H_{2}$ is positive and non-decreasing on $T_{0},T_{1}]$. Then, as shown in [15],
$I^{(1)}]_{t}=[D]_{t}=\int_{T_{0}}^{t}H_{2}(s)\,\mathrm{d}H_{1}(s)-\int_{T_{0}}^{t}H_{1}(s)\,\mathrm{d}H_{2}(s),$ (5.36)
@@
and $H_{1}/H_{2}$ is continuous and monotone, so of bounded variation and differentiable almost everywhere. Recalling that in the standard RAP case we have
@@
$\mathrm{d}M_{t}=\frac{\mathrm{Var}[X_{1}\mid X_{0},I_{t}^{(1)}]}{H_{1}(T_{1})H_{2}(t)-H_{1}(t)H_{2}(T_{1})}\times\left(\left(\frac{Z_{t}(H_{1}^{\prime}(t)H_{2}(T_{1})-H_{1}(T_{1})H_{2}^{\prime}(t))-M_{t}(H_{1}^{\prime}(t)H_{2}(t)-H_{1}(t)H_{2}^{\prime}(t))}{H_{1}(T_{1})H_{2}(t)-H_{1}(t)H_{2}(T_{1})}-J_{t}\right)\,\mathrm{d}t+\,\mathrm{d}I^{(1)}_{t}\right)$,
@@
$\mathrm{d}N_{t}{\mathrel{\mathop{\ordinarycolon}}=}\left(\frac{Z_{t}(H_{1}^{\prime}(t)H_{2}(T_{1})-H_{1}(T_{1})H_{2}^{\prime}(t))-M_{t}(H_{1}^{\prime}(t)H_{2}(t)-H_{1}(t)H_{2}^{\prime}(t))}{H_{1}(T_{1})H_{2}(t)-H_{1}(t)H_{2}(T_{1})}-J_{t}\right)\,\mathrm{d}t+\,\mathrm{d}I^{(1)}_{t}$, (5.40)
@@
$M_{t}=X_{0}+\int_{T_{0}}^{t}\frac{\mathrm{Var}[X_{1}\mid X_{0},I_{t}^{(1)}]}{H_{1}(T_{1})H_{2}(t)-H_{1}(t)H_{2}(T_{1})}\,\mathrm{d}N_{t}.$ (5.41)
As shown below, the process $(N_{t})_{t\in[T_{0},T_{1}]}$, defined by the SDE (5.40) and initial condition $N_{T_{0}}=0$, is a martingale, and can be used to construct a standard Brownian motion on $T_{0},T_{1}]$.
@@
Let $I\subseteq\mathbb{R}^{+}$ be an interval or the real positive line. A stochastic process $(Y_{t})_{t\in I}$ is said to be standard Brownian motion on an interval $a,b]\subseteq I$ if there exists a standard Brownian motion $(B_{t})_{t\geqslant 0}$ such that $Y_{t}\overset{a.s.}{=}B_{t-a}$ for all $t\in[a,b]$.
@@
Under the conditions in Proposition 5.32, the process
$(W_{t})_{t\in[T_{0},T_{1}]}$ defined by
$W_{t}=\int_{T_{0}}^{t}\frac{1}{\sqrt{H_{1}^{\prime}(s)H_{2}(s)-H_{1}(s)H_{2}^{\prime}(s)}}\,\mathrm{d}N_{s}$ (5.42)
is an $(\mathcal{F}_{t}^{I})$-adapted standard Brownian motion on $T_{0},T_{1}]$.